# TNPAR: Topological Neural Poisson Auto-Regressive Model for Learning Granger Causal Structure from Event Sequences

### v1

@@ 5.3 Results on Simulated Data
When compared with the results of PCMCI, ADM4, and MLE_SGL (which do not consider the topological network behind the event sequences), both TNPAR and THP achieve superior performance across all cases. These results suggest that incorporating the topological network assists in learning Granger causality. As can be seen from Fig. 3, TNPAR and its variants significantly outstrip other baselines under most settings, demonstrating the efficacy of the amortized inference technique that treats Granger causal structure as a latent variable.
Across varying parameters $\alpha$, $\mu$, and $\Delta$, the F1 scores of the TNPAR algorithm markedly exceed those of other methods. These results signify that our method is relatively insensitive to the parameters of the generation process, validating the capability of the TNPAR model to effectively model the generation process of event sequences. Under conditions with varying sample sizes, or changing numbers of event types or nodes, the TNPAR algorithm achieves the highest F1 scores, illustrating the relative robustness of the TNPAR algorithm, attributable to its foundation in the neural point process. Even in instances with small sample sizes or high dimensional event types, the TNPAR algorithm still outperforms other methods, corroborating the effectiveness and robustness of TNPAR.
Of all the baseline methods, THP yields the best results since it considers the topological network underlying the event sequences. In comparison to PCMCI, both ADM4 and MLE_SGL methods perform better because they incorporate a sparsity constraint, something PCMCI does not consider.
5.4 Results on Metropolitan Cellular Network Alarm Data
The experimental results of all algorithms on real-world data are presented in Table 1. From these results, it can be seen that the models incorporated the topological network (including TNPAR, TNPAR_NC, THP) outperform that those rely on the i.i.d. assumption. Specifically, the F1 score of TNPAR is higher than that of THP, demonstrating that the TNPAR algorithm can achieve the best trade-off between Precision and Recall without needing a prior distribution assumption. Although the Precision of TNPAR is marginally lower than that of THP, PCMCI, and ADM4, TNPAR exhibits the highest Recall. These results indicate that TNPAR is relatively insensitive to the weak Granger causal strength between event types, which accounts for the slightly higher SHD of TNPAR compared to other baselines. Moreover, TNPAR attains the best SID relative to other methods. These results underscore TNPAR’s superior capability for causal inference Peters and Bühlmann (2015).
@@ 5.4 Results on Metropolitan Cellular Network Alarm Data
Algorithm Precision Recall F1 SID SHD
MLE_SGL 0.39285 0.11702 0.18032 263 100
ADM4 0.49230 0.34042 0.40251 292 95
PCMCI 0.46774 0.30851 0.37179 306 102
THP 0.57894 0.35106 0.43708 266 85
TNPAR_NT 0.30830 0.48271 0.37593 257.00 150.41
TNPAR_NC 0.40893 0.60382 0.48651 271.33 123.49
TNPAR 0.45690 0.64563 0.53509 241.55 105.45
5.5 Ablation Study
@@
Ablation experiments were conducted on both simulated and real-world data to assess the effectiveness of incorporating topological information and regularization in TNPAR. The experimental results generally reveal a significant improvement in the performance of TNPAR compared to both TNPAR_NT and TNPAR_NC. These results underscore the importance of incorporating topological information, as well as acyclic and sparse regularization. Additionally, in most instances, both TNPAR_NT and TNPAR_NC perform comparably to THP, despite the fact that the THP fully leverages topological information and sparsity constraints. These results demonstrate the robustness of the TNPAR algorithm.
5.6 Case Study
@@
In the real-world data experiment, our method successfully infers the Granger causality across the nodes of the topological network, aligning with experts’ knowledge. From the results, we observe that the count of causal edges diminishes as the geodesic distance, denoted as $k$, increases. This finding aligns with our intuition that certain events are influenced solely by their node. For instance, the causal relationship $BD\_STATUS\rightarrow TU\_AIS$ is only valid for $k=0$. Here, $BD\_STATUS$ represents a device disconnecting from a node, and $TU\_AIS$ signifies that some tasks running on this node have been disrupted. Another notable discovery is that certain edges exclusively transit across nodes, suggesting that these event types are triggered only by their neighbors. Consider, for example, consider $ETH\_LINK\_DOWN\rightarrow MW\_LOF$, where $ETH\_LINK\_DOWN$ denotes an error on the Ethernet interface in a downstream node, and $MW\_LOF$ indicates disconnection of an upstream node.
6 Conclusion
In this paper, we present a Granger causal discovery algorithm for topological event sequences. Harnessing the pre-existing topological network and regarding the causal structure as a latent variable, we devise a method to tackle the non-i.i.d. challenge in causal discovery. This approach also addresses the complexity of unified modeling of the topological network and latent causal structure. Specifically, we first propose a comprehensive topological neural Poisson auto-regressive model to capture the generation process of topological event sequences. Subsequently, we employ amortized inference to infer the latent Granger causal structure. Lastly, we integrate these two stages into a unified likelihood function that can be optimized through an end-to-end solution. Experimental results from both simulation and real-world data corroborate the efficacy of our proposed method.

### v2

@@ Introduction
As shown in Fig. 3, when compared with the results of PCMCI, ACD, ADM4, and MLE_SGL (which do not consider the topological network behind the event sequences), both TNPAR and THP achieve superior performance across all cases. These results suggest that incorporating the topological network assists in learning Granger causality. Moreover, TNPAR and its variants significantly outstrip other baselines under most cases, demonstrating that the neural point process approach is powerful for modeling the generation process and the amortized inference technique is effective for Granger causal discovery.
Traversing the value of parameters $\alpha$, $\mu$, and $\Delta$, the F1 scores of the TNPAR are generally better than those of other methods. These results signify that our method is relatively insensitive to the parameters of the generation process. Under settings with varying sample sizes, or changing numbers of event types or nodes, TNPAR achieves the highest F1 scores, illustrating the robustness of TNPAR. Even in cases with small sample sizes or high dimensional event types, TNPAR still outperforms other methods, corroborating the effectiveness of TNPAR.
Among the baseline methods, THP delivers the best results as it considers the topological network. Both ADM4 and MLE_SGL outperform PCMCI and ACD, which demonstrates the efficacy of the point process methods in causal discovery for event sequences.
Results on Metropolitan Cellular Network Alarm Data
The experimental results of all algorithms on real-world data are presented in Table 1. From these results, it can be seen that the models incorporated the topological network (including TNPAR, TNPAR_NC, THP) outperform that those rely on the i.i.d. assumption. Specifically, the F1 score of TNPAR is higher than that of THP, demonstrating that TNPAR can achieve the best trade-off between Precision and Recall without needing a prior distribution assumption. Although the Precision of TNPAR is marginally lower than that of THP, PCMCI, and ADM4, TNPAR exhibits the highest Recall. These results indicate that TNPAR is relatively insensitive to the weak Granger causal strength between event types, which accounts for the slightly higher SHD of TNPAR compared to other baselines. Moreover, TNPAR attains the best SID relative to other methods except ACD. These results underscore TNPAR’s superior capability for causal inference (Peters and Bühlmann 2015).
Algorithm Prec. Rec. F1 SID SHD
ACD 0.25 0.32 0.27 230.0 127.0
MLE_SGL 0.39 0.12 0.18 263.0 100.0
ADM4 0.49 0.34 0.40 292.0 95.0
PCMCI 0.47 0.31 0.37 306.0 102.0
THP 0.58 0.35 0.44 266.0 85.0
TNPAR_NT 0.31 0.48 0.38 257.0 150.4
TNPAR_NC 0.41 0.60 0.49 271.3 123.5
TNPAR_MG 0.35 0.31 0.33 272.5 115.2
TNPAR 0.46 0.65 0.54 241.6 105.5
@@
Ablation Study
@@
Ablation experiments were conducted on both simulated and real-world data to assess the effectiveness of incorporating topological information and regularization in TNPAR. The experimental results given in Fig. 3 reveal a significant improvement in the performance of TNPAR compared to TNPAR_NT, TNPAT_MG, and TNPAR_NC. These results underscore the importance of incorporating topological information, as well as acyclicity and sparsity regularization. Additionally, in most instances, TNPAR_NT, TNPAR_MG, and TNPAR_NC perform comparably to THP, despite the fact that the THP fully leverages topological information and sparsity constraint. These results demonstrate the robustness of TNPAR.
Case Study
@@
In the real-world data experiment, our method successfully infers the Granger causality across the nodes of the topological network, aligning with experts’ knowledge. From the results, we observe that the count of causal edges diminishes as the geodesic distance, denoted as $k$, increases. This finding aligns with our intuition that some events are influenced solely by their node. For instance, the causal relationship $BD\_STATUS\rightarrow TU\_AIS$ is only valid for $k=0$. Here, $BD\_STATUS$ represents a device disconnecting from a node, and $TU\_AIS$ signifies that some tasks running on this node have been disrupted. Another notable discovery is that some edges exclusively transit across nodes, suggesting that these event types are triggered only by their neighbors. For example, consider $ETH\_LINK\_DOWN\rightarrow MW\_LOF$, where $ETH\_LINK\_DOWN$ denotes an error on the Ethernet interface in a downstream node, and $MW\_LOF$ indicates disconnection of an upstream node.
Conclusion
In this paper, we studied the Granger causal discovery problem on the topological event sequences. By leveraging the prior topological network and regarding the causal structure as a latent variable, we successfully addressed the challenge of unified modeling the topological network and latent causal structure, which enable us to propose the amortized causal discovery method. The experimental results on both simulation and real-world data demonstrate the effectiveness of our proposed method. TNPAR offers a viable solution for causal discovery in real-world settings with non-i.i.d. data. To further enhance its effectiveness, our future work will focus on improving sample efficiency, integrating prior knowledge, and so on.
Acknowledgments
This research was supported in part by National Key R&D Program of China (2021ZD0111501), National Science Fund for Excellent Young Scholars (62122022), Natural Science Foundation of China (61876043, 61976052, 62206061, 62206064), the major key project of PCL (PCL2021A12), Guangzhou Basic and Applied Basic Research Foundation (2023A04J1700).
Supplementary Material
In this supplementary material, we summarize the methods related to our work in Appendix A. Then we provide the proof for Proposition 1 in Appendix B. After that, Appendix C describes the details of the baseline methods in experiments. At last, more results in simulated data are given in Appendix D.
A. Related Work Summarization
In this section, we explicitly compare our proposed method with the related methods from four perspectives in Table A.2. From this table, we can find that our method is the only method for non-i.i.d. topological event sequences.
Table A.2: Related methods for causal discovery.
Methods Categories Density Estimator Probabilistic Relation Non-i.i.d.
MLE_SGL Hawkes process-based Gaussian-based kernel No No
ADM4 Hawkes process-based exponential-based kernel No No
THP Hawkes process-based exponential-based kernel No Yes
PCMCI constraint-based - No No
ACD auto-regressive-based - Yes Yes
TNPAR neural point process-based neural network Yes Yes
B. Proof for Proposition 1
Proposition 2.
Given the Granger causal structure $\mathcal{G}_{V}(\mathbf{V},\mathbf{E}_{V})$, the topological network $\mathcal{G}_{N}(\mathbf{N},\mathbf{E}_{N})$ and the max geodesic distance $K$, along with the assumption that the data generation process adheres to Eq. (4), we can deduce that $v_{i}\rightarrow v_{j}\notin\mathbf{E}_{V}$, if and only if $A_{k}^{v_{i},v_{j}}=0$ for every $k\in\{0,...,K\}$.
Proof.
Let $O_{t}^{v_{j},n_{k}}$ and $\hat{O}_{t}^{v_{j},n_{k}}$ represent the outputs for $\mathrm{g}(\mathbf{O}^{v_{1},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{i},n_{1}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{i},n_{|N|}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|},n_{|N|}}_{t-1:t-\Omega})$ and $\mathrm{g}(\mathbf{O}^{v_{1},n_{1}}_{t-1:t-\Omega},...,\mathbf{\hat{O}}^{v_{i},n_{1}}_{t-1:t-\Omega},...,\mathbf{\hat{O}}^{v_{i},n_{|N|}}_{t-1:t-\Omega},...,\mathbf{O}^{v_{|V|},n_{|N|}}_{t-1:t-\Omega})$ in Definition 3, respectively.
$\Longrightarrow$: If $v_{i}\rightarrow v_{j}\notin\mathbf{E}_{V}$, then according to Definition 3 and Eq. (4), it necessitates that $O_{t}^{v_{j},n_{k}}=\hat{O}_{t}^{v_{j},n_{k}}$. Conversely, according to Eq. (4), if $A_{k}^{v_{i},v_{j}}\neq 0$ for any $k\in\{0,...,K\}$, the inputs for $\mathrm{g}(\cdot)$ will differ, potentially altering its output, (i.e., $O_{t}^{v_{j},n_{k}}\neq\hat{O}_{t}^{v_{j},n_{k}}$), leading to a contradiction. Hence, if $v_{i}\rightarrow v_{j}\notin\mathbf{E}_{V}$, it logically follows that $A_{k}^{v_{i},v_{j}}=0$ for every $k\in\{0,...,K\}$ must hold.
[…]
C. Details of the Methods in Experiments
This section provides the details and parametric settings of methods chosen by cross-validation in the experiments. We first begin with the baseline methods as follows:
  • PCMCI: The algorithm uses the independence test in the first stage to remove the irrelevant edges between event types at different time lags from a complete undirected graph. Subsequently, the momentary conditional independence (MCI) test is introduced in the second stage to determine the arrows for the undirected edges. In these experiments, we set the time interval to 5 seconds, the max time lag to 2, and the significant level of independence test to 0.01. Because the input data is discrete, the conditional mutual information (CMI) test is used as the independence test method.
  • MLE_SGL: The algorithm combines a maximum likelihood estimator (MLE) with a sparse group lasso (SGL) regularizer for the Hawkes process model to learn the Granger causal structure between event types. In detail, MLE_SGL uses a linear combination of Gaussian base functions to implement the intensity function in the Hawkes process. In these experiments, the number of the Gaussian base function is set to 10 and the coefficient of the group lasso regularization is set to 0.5.
  • ACD: The algorithm introduces an amortized model that infers causal relations across time series with different underlying causal graphs, which share relevant information. Specifically, ACD encompasses an amortized encoder that predicts the edges in the causal graph, and a decoder models the dynamics of the system under the predicted causal relations. In these experiments, we aggregate the event sequences into time series using a window size of 5 seconds. For the algorithm’s execution, we employ the original MLP encoder and decoder, utilizing the default parameters given by the author.
  • ADM4: The algorithm learns the Granger causal structure between event types by combining techniques of alternating direction method of multipliers and majorization minimization for Hawkes process. Different from MLE_SGL, it uses the exponential kernel to implement the intensity function with a decay parameter $\beta$. To impose priors on sparsity and low-rank causal structure, both the nuclear norm and $\ell_{1}$-norm regularization are adopted. In our experiment, the decay $\beta$ is set to 0.4 and the coefficients of $\ell_{1}$-norm and nuclear regularization are set to 0.05.
  • THP: The algorithm extends the time-domain convolution property of the Hawkes process to the time-graph domain for learning causal structure among event types generated in a topological network. In these experiments, we follow the original implementation of THP and use an exponential kernel to model its intensity function. The decay rate $\delta$ of the exponential kernel is set to 0.1.
To evaluate the modules of the proposed method, we devise three variants of our method: TNPAR_NT, TNPAR_MG, and TNPAR_NC. The details of our method and its variants in the experiments are as follows.
  • TNPAR: In this work, we propose TNPAR for Granger causal discovery on topological event sequences. By incorporating the topological network $\mathbf{B}_{0:K}$ and the casual structure $\mathbf{A}_{0:K}$, TNPAR solves the problem of causal discovery from non-i.i.d. data. In addition, TNPAR is applied to optimize the parameters with a sparsity regularization $\mathcal{L}_{s}$ and an acyclicity regularization $\mathcal{L}_{c}$ on $\mathbf{A}_{0:K}$. In these experiments, to obtain a sparse and acyclic causal structure, $\mathcal{L}_{s}$ is set to 1e-4 and $\mathcal{L}_{c}$ is 1e-10.
  • TNPAR_NT: To evaluate the effectiveness of introducing topological neighbors’ information into TNPAR, TNPAR_NT only uses the historical information on its own node, without considering the topological information. In our experiment, $\lambda_{s}$ is set to 1e-4 and $\lambda_{c}$ is set to 1e-10.
  • TNPAR_MG: To evaluate the effectiveness of introducing the topological network into TNPAR, we merge all event sequences produced by the topological nodes, which is used as the training data for TNPAR_MG. In our experiment, $\lambda_{s}$ is set to 1e-9 and $\lambda_{c}$ is set to 1e-7.
  • TNPAR_NC: To evaluate the effectiveness of the acyclic and sparse regularization in TNPAR, TNPAR_NC removes the acyclicity and sparsity regularization from the loss function. That is, $\lambda_{s}$ and $\lambda_{c}$ are both set to 0.
D. Details of the Results in Simulation Dataset
In this section, we further provide F1, Precision, Recall, SID, and SHD of baseline methods and our method in each experimental setting. The results are given in Fig. D.4 - D.8. All experiments are conducted on 60 cores AMD EPYC 7543 Processor and 4 * RTX 3090 with 320GB RAM. The operating system is Linux. In addition, the relevant software libraries and frameworks include the PyTorch 1.7.0 version, Python 3.8 version, and Cuda 11.0.
From the experimental results, we can find that TNPAR achieves the best performance compared to all baselines (MLE_SGL  ADM4, PCMCI, and THP), which demonstrates the superiority of TNPAR. In addition, with the varying parameters $\alpha$, $\mu$, and $\Delta$ in the data generation process, the performance of TNPAR is more robust than that of the baselines in all cases. The reason is that TNPAR is implemented by the neural point process, which holds a powerful capacity in modeling the generation process. Moreover, even in the cases of low sample size, high event number, or high node number, the experimental results of TNPAR remain well, which demonstrates its excellence.
Compared with the variants (TNPAR_NT, TNPAR_MG, and TNPAR_NC), TNPAR achieves the best F1 score and SHD in all ablation studies, which proves the necessity of introducing the topological network, sparse constraint, and acyclic constraint. The Recall of TNPAR_NT and TNPAR_MG are lower than that of TNPAR because they do not learn the causal relationships across nodes while TNPAR does. In addition, the Precision of TNPAR_NC is lower than that of TNPAR, which means there are more redundant edges in the result of TNPAR_NC than TNPAR due to the lack of constraints.
(a) Sensitivity to Range of $\alpha(\times 1e-2)$
(b) Sensitivity to Range of $\mu(\times 1e-5)$
(c) Sensitivity to Sample Size
(d) Sensitivity to $\Delta$
(e) Sensitivity to Num. of Event Types
(f) Sensitivity to Num. of Nodes
Figure D.4: F1 scores on the simulated data
(a) Sensitivity to Range of $\alpha(\times 1e-2)$
(b) Sensitivity to Range of $\mu(\times 1e-5)$
(c) Sensitivity to Sample Size
(d) Sensitivity to $\Delta$
(e) Sensitivity to Num. of Event Types
(f) Sensitivity to Num. of Nodes
Figure D.5: Precision on the simulated data
(a) Sensitivity to Range of $\alpha(\times 1e-2)$
(b) Sensitivity to Range of $\mu(\times 1e-5)$
(c) Sensitivity to Sample Size
(d) Sensitivity to $\Delta$
(e) Sensitivity to Num. of Event Types
(f) Sensitivity to Num. of Nodes
Figure D.6: Recall on the simulated data
(a) Sensitivity to Range of $\alpha(\times 1e-2)$
(b) Sensitivity to Range of $\mu(\times 1e-5)$
(c) Sensitivity to Sample Size
(d) Sensitivity to $\Delta$
(e) Sensitivity to Num. of Event Types
(f) Sensitivity to Num. of Nodes
Figure D.7: SID on the simulated data
(a) Sensitivity to Range of $\alpha(\times 1e-2)$
(b) Sensitivity to Range of $\mu(\times 1e-5)$
(c) Sensitivity to Sample Size
(d) Sensitivity to $\Delta$
(e) Sensitivity to Num. of Event Types
(f) Sensitivity to Num. of Nodes
Figure D.8: SHD on the simulated data